OpenStudy (anonymous):

Find the cube roots of 27(cos 330° + i sin 330°).

OpenStudy (anonymous):

Lol....u Know De-Morvies Theorem

OpenStudy (anonymous):

I think it is (costheta + i sintheta)^n = (cosn theta + i sin n theta)

OpenStudy (anonymous):

(cosn theta + i sin n theta) = e^(i theta)

OpenStudy (anonymous):

27(cos 330° + i sin 330°) = 27 e^(i 11pie/6)

OpenStudy (anonymous):

now find its Cube Root....)

OpenStudy (anonymous):

I think I did this wrong. 44796.22502i?

OpenStudy (anonymous):

@Yahoo!

OpenStudy (anonymous):

Your Euclids theorem is wrong I think.. @Yahoo!

OpenStudy (anonymous):

\[\large \cos(\theta) + i \sin(\theta) = e^{i \theta}\] According to Euler Therorem..

OpenStudy (anonymous):

This can be solved by using De-Moivre's theorem only..

OpenStudy (anonymous):

Tell me, according to that theorem, what will you get after this?? \[\large (27(\cos(330) + i \sin(330))^{\frac{1}{3}} = ??\]

OpenStudy (anonymous):

oh...Sorry....i understood this question Differently..

OpenStudy (anonymous):

So what do i do?

OpenStudy (anonymous):

Use here De-Moivre's theorem.. Here n = 1/3..

OpenStudy (anonymous):

I think you have written formula for the theorem above.. Use that @Itel

Parth (parthkohli):

@Yahoo! @waterineyes It's Euler's Formula :p

OpenStudy (anonymous):

More precisely, that is one of Euler's Identity.. @parthkohli ..

OpenStudy (anonymous):

I keep getting 26.58980933-4.688500797i

zepdrix (zepdrix):

\[\huge 27(\cos 330^o+i \sin 330^o)=27e^{\frac{11\pi}{6}i}\]Taking the cube root gives us,\[\huge \left(27e^{\frac{11\pi}{6}i}\right)^{1/3}\quad=\quad 3e^{\frac{11\pi}{18}i}\]This angle converted to degrees is,\[\huge \frac{11\pi}{18}\cdot \frac{180}{\pi} \quad = \quad 110^o\]Yahoo had the right idea. You can certainly do it this way if you want :O Using that theorem is probably faster though :)

zepdrix (zepdrix):

\[\large 3(\cos 110^o+i \sin 110^o)\]

OpenStudy (anonymous):

It keeps popping up as [Math processing Error] in your response. So I can't really figure out what it is saying. Sorry. I'm just completely confused on this one. I tried using the theorem and I got this way off the wall answer. I tried some other relational thing I found on google and that gave me an even more bizarre answer.

OpenStudy (anonymous):

Hey when you get [Math processing Error] then reload or refresh your page.. You will see everything then..

OpenStudy (anonymous):

\[\large [27(\cos(330) + i \cdot \sin(330))]^{\frac{1}{3}} = (27)^{\frac{1}{3}}(\cos(\frac{1}{3} \times 330) + i \cdot \sin(\frac{1}{3} \times 330))\]

Parth (parthkohli):

@waterineyes Let me get it straight :) \[\textbf{EULER'S IDENTITY} \ \ \ \ e^{\pi i} + 1 = 0\]\[\textbf{EULER'S FORMULA} \ \ \ \ e^{i \phi} = \cos(\phi) + i\sin(\phi)\]

OpenStudy (anonymous):

Finally you will get as @zepdrix said: \[\implies \color{blue}{3 [\cos(110) + i \cdot \sin(110)]}\]

OpenStudy (anonymous):

Euler gave two Identities : \[e^{i \theta} = \cos(\theta) + i \sin(\theta)\] \[e^{-i \theta} = \cos(\theta) - i \sin(\theta)\] If I am not wrong then..

Parth (parthkohli):

Yes, the second is just a derivative of the second. :)

OpenStudy (anonymous):

What??

OpenStudy (anonymous):

Second is just derivative of second?? Reflexive Property?? Ha ha ha..

Parth (parthkohli):

Dusra wala formula pehle se derive hua hai kaakey :D

Parth (parthkohli):

It's pretty easy to do so:\[e^{-i\phi} = e^{i\times(-\phi)} = \cos(-\phi) + i\sin(-\phi) = \cos(\phi) - i\sin(\phi). \]Knowing that \(\sin(-x) = -\sin(x)\) and \(\cos(-x) = \cos(x).\)

OpenStudy (anonymous):

Yes, the second is just a derivative of the second. :) The second time you have written second should be first I think.. :P

Parth (parthkohli):

The second is just a derivative of the first*

OpenStudy (anonymous):

Okay.. Yes you are right.. and thanks for clearing me straight on Euler's Identity and Formula..

zepdrix (zepdrix):

It's not the derivative of the first. It's just the identity for -theta. Given that Cosine is even, Sine is an odd function, the negatives move around nicely. If it was the derivative it would look something like...\[\large ie^{i \theta}=-\sin \theta+i \cos \theta\] I'm not sure what that formula would tell us :D heh

OpenStudy (anonymous):

Parthkohli, just meant to say that second one can be easily derived from First formula.. He is not saying about Differentiation or Derivative..

zepdrix (zepdrix):

Oh my mistake ^^